# chebgreen—: Learning and Interpolating Continuous Empirical Green’s Functions from Data

### v2

@@ 2.4 Approximating a bivariate function, $g(x,y)$, with chebfun2
Approximating a function with a chebfun allows for constructing infinite-dimensional analogues of matrices, which are called Quasimatrices. A column Quasimatrix can be thought of as ×$\infty\times K$ matrices. In other words, they have finitely many columns, each column being represented by a square-integrable function, in this case a chebfun.
@@
Consider a ×$\infty\times K$ column Quasimatrix $\mathcal{A}$, where $\{\mathcal{A}_{k}\}_{k=1}^{K}$ are chebfun-s on some domain $\Omega\subset\mathbb{R}$.
@@
In this paper, we will denote the adjoint of a column Quasimatrix $\mathcal{A}$ as $\mathcal{A}^{*}$, and note that, given another column Quasimatrix $\mathcal{C}$, the inner product between two quasimatrices (an outer product on the space $(L^{2}(\Omega))^{K}$) is given by [40]:
@@
The idea of employing a polynomial approximation of a univariate function on a Chebyshev grid can be extended to approximate bivariate functions. chebfun2 [32] constructs approximate $g(x,y)\approx C(y)DR(x)$ [^1] This is analogous to a CUR-decomposition for a matrix but we use $D$ instead of $U$ to avoid confusion with our terminology for left singular vectors. It constructs a rank $K$ approximant as follows:
@@
The pivots are chosen at the locations where the absolute error between the current approximant and the bivariate function $G(x,y)$ is maximum. Repeating this procedure iteratively for $K$ steps, chebfun2 constructs an approximation for $g(x,y)$. A more detailed overview of the workings of chebfun2 can be found in the original work [32].
@@ 2.5 Constructing a singular value expansion of $g(x,y)$
Constructing this Singular Value Expansion thus requires computing a QR decomposition of a Quasimatrix. The details of the same can be found in Trefethen’s work on Householder triangularization of a Quasimatrix [41]. The main details are summarized here.
[…]
@@
$\bm{E}$ will be an ×$\infty\times n$ quasimatrix , fixed in advanced, with orthonormal columns $\{e_{j}\}$, and $\bm{S}$ is an $n\times n$ sign matrix, so as to ensure the diagonal of $\bm{R}$ is real and non-negative. In practice, $e_{j}$ is taken to be a multiple of the $(j-1)$th Legendre polynomial $P_{j-1}(x)$, scaled to $[a,b]$. The Householder reflector $\bm{H}_{k}$ is a self-adjoint operator acting on $L^{2}[a,b]$, chosen so as to map a certain function to another function of equal norm in the space spanned by $e_{1},\dots,e_{k}$.
@@
The process of constructing an SVE with chebfun2 [32] is outlined in Algorithm 1.
@@ 2.7 Interpolating the solution on a manifold
The algorithm proposed by Praveen et al.[30] relies on the method of interpolating in a tangent space of the compact Stiefel manifold. The compact $k-$Stiefel manifold of $\mathbb{R}^{m}$ is defined to be the manifold of equivalence classes of $m\times k$ matrices with orthonormal columns, where $U$ and $U^{\prime}$ are consiedered equivalent when their columns span the same subspace of $\mathbb{R}^{m}$ [42]. In other words, we may think of it as
@@
Their previous algorithm takes the SVDs of known Green’s functions, $G_{i}=\Phi_{i}\Sigma_{i}\Phi_{i}^{T}$, at parameters $\theta_{i}$, chooses an index $j$ that is closest to the unseen parameter, $\theta^{*},$ and lets $\Phi_{j}=\hat{\Phi}$ be the reference basis (base point) for the interpolation. The other known orthonormal matrices, $\Phi_{i}$, represented as elements on the Stiefel manifold, are then mapped to the tangent space at $\hat{\Phi}$ via an approximation to the logarithmic map [43]. These points in the tangent space are then interpolated via Lagrange polynomial interpolation, and the matrix, $\Phi^{*}$, for the unseen parameter, $\theta^{*}$, is obtained by mapping the interpolated point back to the Stiefel manifold from the tangent space using an exponential mapping [43].
@@
In order to generalize the manifold interpolation, we first need to find a space analogous to the finite-dimensional Stiefel manifold, and to identify its tangent space [40]. We write elements, $\Phi$, of $H$ in row-vector notation as $\Phi=(\Phi_{1},...,\Phi_{k}),$ so that we can multiply quasimatrices (viewed here as vectors in $H$) on the right by a $k\times k$ matrix to get another vector in $H.$ The space $H$ admits an outer product [40]
$\Phi^{*}\Psi:=\left[\begin{array}[]{ccc}\left(\Phi_{1},\Psi_{1}\right)_{L^{2}(\Omega)}&\cdots&\left(\Phi_{1},\Psi_{k}\right)_{L^{2}(\Omega)}\\ \vdots&\ddots&\vdots\\ \left(\Phi_{k},\Psi_{1}\right)_{L^{2}(\Omega)}&\cdots&\left(\Phi_{k},\Psi_{k}\right)_{L^{2}(\Omega)}\end{array}\right]\in\mathbb{R}^{k\times k}$ (12)
@@
$(\Phi,\Psi)_{H}:=\sum_{j=1}^{k}\left(\Phi_{j},\Psi_{j}\right)_{L^{2}(\Omega)}=\operatorname{tr}(\Phi^{*}\Psi).$ (13)
Here, we use the suggestive notation $\Phi^{*}\Psi$ for the outer product on $H$ to emphasize that this is both distinct from, and analogous with, the matrix transpose [44]. We then define
@@
The inner product on $H$ induces a Riemannian metric, and turns $S_{k}(H)$ into a Riemannian submanifold of $H$ which we will call the Stiefel manifold, and will denote from here on simply by $S_{k}$ [40]. It can be shown that its tangent space at a point $\Phi$ is
@@
This projection mapping is our infinite dimensional analogue to Step 4 Part 1 of the previous algorithm [30].
In order to extend the third part of Step 4 of that same algorithm to the Quasimatrix generalization, we need to establish a valid approximation to the exponential map so that we can map interpolated points in the tangent space back to the Stiefel manifold. The key concept is that of a retraction on a manifold, which may be thought of as a first order approximation to the exponential map. In their previous algorithm Praveen et al.[30] use QR factorization as their approximation to the exponential map, as it is retraction on the finite dimensional Stiefel manifold. Altman et al. [40] show that Quasimatrix QR factorization is a retraction on our infinite dimensional Stiefel manifold $S_{k}$, so it follows that we may simply use the Quasimatrix version of QR factorization as our approximation to the exponential map in our updated algorithm.
@@
    $\Gamma^{\mathcal{U}_{0}}_{j}$ $=\mathcal{U}(s;\theta_{j}))-\mathcal{U}(s;\theta_{0}))\ \text{sym}\left(\mathcal{U}^{*}(s;\theta_{0}))\mathcal{U}(s;\theta_{j}))\right),$
    $\Gamma^{\mathcal{V}_{0}}_{j}$ $=\mathcal{V}(x;\theta_{j}))-\mathcal{V}(x;\theta_{0}))\ \text{sym}\left(\mathcal{V}^{*}(x;\theta_{0}))\mathcal{V}(x;\theta_{j}))\right),\quad\text{sym}\left(Y\right)\coloneqq(Y+Y^{*})/2.$
@@
2.7.1 Correcting eigenmodes sign and order
@@
We observe that the first two singular functions swap at some critical value of $\theta$. The $k^{\text{th}}$ singular value for the Green’s functions of the operator $\mathcal{L}_{\theta}$ are given by $\sigma_{k}=1/(\theta^{2}-(k\pi)^{2})$. If we compare the first two modes, they satisfy $|\sigma_{1}|>|\sigma_{2}|$, when $\theta<\sqrt{5/2}\pi$, and $|\sigma_{1}|\leq|\sigma_{2}|$ otherwise. This implies that the first two eigenmodes are swapped when $\theta>\sqrt{5/2}\pi$, as illustrated by Fig. 7, and that the manifold interpolation technique will perform poorly in such cases. The same phenomena occurs for the higher eigenmodes at larger values of $\theta$. We propose to reorder the singular functions and the associated singular values of the discovered basis at the given interpolation parameter to match the ones at the origin point $\theta_{0}$, where we lift to the tangent space of the infinite-dimensional analogue for Stiefel manifold. For a given parameter, $\theta_{j}$, and mode number, $1\leq k\leq K$, we select the $k^{\text{th}}$ left singular function at $\theta_{j}$ to be the one with minimal angle with the $k^{\text{th}}$ left singular function at parameter $\theta_{0}$, i.e.,
$\mathcal{U}_{k}(y;\theta_{j})=\mathcal{U}_{\ell^{\prime}}(y;\theta_{j}),\quad\text{where}\quad\ell^{\prime}=\text{match}(k),$
$\text{match}(k)=$ $\operatorname*{arg\,max}_{1\leq\ell\leq K,\ell\neq\text{match}(k^{\prime})\forall 1\leq k^{\prime}<k}|\langle\mathcal{U}_{\ell}(y;\theta_{j}),\mathcal{U}_{k}(y;\theta_{0})\rangle|,\quad 1\leq k\leq K.$
@@
[…]
Figure 7: (A) Magnitude of the first two eigenvalues of the Green’s function associated with the one-dimensional Helmholtz equation. The first two corresponding eigenmodes are swapped when the frequency, $\theta$, increases beyond the critical value of $\theta_{\text{crit}}=\sqrt{5/2}\pi\approx 5$, as displayed in (B) and (C).

### v5

@@ 2.4 Approximating a bivariate function, $g(x,y)$, with chebfun2
Approximating a function with a chebfun allows for constructing infinite-dimensional analogues of matrices, which are called Quasimatrices. A column Quasimatrix can be thought of as $\infty\times K$ matrices. In other words, they have finitely many columns, each column being represented by a square-integrable function, in this case a chebfun.
@@
Consider a $\infty\times K$ column Quasimatrix $\mathcal{A}$, where $\{\mathcal{A}_{k}\}_{k=1}^{K}$ are chebfun-s on some domain $\Omega\subset\mathbb{R}$.
@@
In this paper, we will denote the adjoint of a column Quasimatrix $\mathcal{A}$ as $\mathcal{A}^{*}$, and note that, given another column Quasimatrix $\mathcal{C}$, the inner product between two quasimatrices (an outer product on the space $(L^{2}(\Omega))^{K}$) is given by [Altmann2022_Stiefel]:
@@
The idea of employing a polynomial approximation of a univariate function on a Chebyshev grid can be extended to approximate bivariate functions. chebfun2 [Townsend_Chebfun2] constructs approximate $g(x,y)\approx C(y)DR(x)$ [^1] This is analogous to a CUR-decomposition for a matrix but we use $D$ instead of $U$ to avoid confusion with our terminology for left singular vectors. It constructs a rank $K$ approximant as follows:
@@
The pivots are chosen at the locations where the absolute error between the current approximant and the bivariate function $G(x,y)$ is maximum. Repeating this procedure iteratively for $K$ steps, chebfun2 constructs an approximation for $g(x,y)$. A more detailed overview of the workings of chebfun2 can be found in the original work [Townsend_Chebfun2].
@@ 2.5 Constructing a singular value expansion of $g(x,y)$
Constructing this Singular Value Expansion thus requires computing a QR decomposition of a Quasimatrix. The details of the same can be found in Trefethen’s work on Householder triangularization of a Quasimatrix [Trefethen2010_QR]. The main details are summarized here.
Let’s assume we want to find a QR decomposition for a $\infty\times n$ Quasimatrix $\bm{A}$, the columns of which are functions of $x\in[a,b]$ ($\bm{A}$ is a linear map from $\mathbb{C}^{n}$ to $L^{2}[a,b]$). Let the notation $\bm{v}^{*}\bm{w}$ denote the inner product and $\|\cdot\|$ be the $L^{2}$ norm on $[a,b]$. We would like to decompose this Quasimatrix as $\bm{A}=\bm{Q}\bm{R}$, such that $\bm{Q}$ is an orthonormal Quasimatrix and $\bm{R}$ is an upper triangular matrix. The aim is to compute this by the numerically more stable method of Householder reflections. This is done by applying a self-adjoint operator, $\bm{H}$, so that one can get the form as shown in Fig. 5. Here, the vertical lines denote that the column is a chebfun.
@@
$\bm{E}$ will be an $\infty\times n$ quasimatrix , fixed in advanced, with orthonormal columns $\{e_{j}\}$, and $\bm{S}$ is an $n\times n$ sign matrix, so as to ensure the diagonal of $\bm{R}$ is real and non-negative. In practice, $e_{j}$ is taken to be a multiple of the (j1)(j-1)th Legendre polynomial $P_{j-1}(x)$, scaled to $[a,b]$. The Householder reflector $\bm{H}_{k}$ is a self-adjoint operator acting on $L^{2}[a,b]$, chosen so as to map a certain function to another function of equal norm in the space spanned by $e_{1},\dots,e_{k}$.
@@
The process of constructing an SVE with chebfun2 [Townsend_Chebfun2] is outlined in Algorithm 1.
@@ 2.7 Interpolating the solution on a manifold
The algorithm proposed by Praveen et al.[Praveen2023_EGF] relies on the method of interpolating in a tangent space of the compact Stiefel manifold. The compact $k$-Stiefel manifold of $\mathbb{R}^{m}$ is defined to be the manifold of equivalence classes of $m\times k$ matrices with orthonormal columns, where $U$ and $U^{\prime}$ are consiedered equivalent when their columns span the same subspace of $\mathbb{R}^{m}$ [Absil2008_Manifolds]. In other words, we may think of it as
@@
Their previous algorithm takes the SVDs of known Green’s functions, $G_{i}=\Phi_{i}\Sigma_{i}\Phi_{i}^{T}$, at parameters $\theta_{i}$, chooses an index $j$ that is closest to the unseen parameter, $\theta^{*},$ and lets $\Phi_{j}=\hat{\Phi}$ be the reference basis (base point) for the interpolation. The other known orthonormal matrices, $\Phi_{i}$, represented as elements on the Stiefel manifold, are then mapped to the tangent space at $\hat{\Phi}$ via an approximation to the logarithmic map [Sternfels2013_Interp]. These points in the tangent space are then interpolated via Lagrange polynomial interpolation, and the matrix, $\Phi^{*}$, for the unseen parameter, $\theta^{*}$, is obtained by mapping the interpolated point back to the Stiefel manifold from the tangent space using an exponential mapping [Sternfels2013_Interp].
@@
In order to generalize the manifold interpolation, we first need to find a space analogous to the finite-dimensional Stiefel manifold, and to identify its tangent space [Altmann2022_Stiefel]. We write elements, $\Phi$, of $H$ in row-vector notation as $\Phi=(\Phi_{1},...,\Phi_{K}),$ so that we can multiply quasimatrices (viewed here as vectors in $H$) on the right by a $K\times K$ matrix to get another vector in $H.$ The space $H$ admits an outer product [Altmann2022_Stiefel]
$\Phi^{*}\Psi=\left[\begin{array}[]{ccc}\left(\Phi_{1},\Psi_{1}\right)_{L^{2}(\Omega)}&\cdots&\left(\Phi_{1},\Psi_{K}\right)_{L^{2}(\Omega)}\\ \vdots&\ddots&\vdots\\ \left(\Phi_{K},\Psi_{1}\right)_{L^{2}(\Omega)}&\cdots&\left(\Phi_{K},\Psi_{K}\right)_{L^{2}(\Omega)}\end{array}\right]\in\mathbb{R}^{K\times K}$ (12)
@@
$(\Phi,\Psi)_{H}=\sum_{j=1}^{K}\left(\Phi_{j},\Psi_{j}\right)_{L^{2}(\Omega)}=\operatorname{tr}(\Phi^{*}\Psi).$ (13)
Here, we use the suggestive notation $\Phi^{*}\Psi$ for the outer product on $H$ to emphasize that this is both distinct from, and analogous with, the matrix transpose [Harms2012_Manifolds]. We then define
@@
The inner product on $H$ induces a Riemannian metric, and turns $S_{K}(H)$ into a Riemannian submanifold of $H$ which we will call the Stiefel manifold, and will denote from here on simply by $S_{K}$ [Altmann2022_Stiefel]. It can be shown that its tangent space at a point $\Phi$ is
@@
This projection mapping is our infinite dimensional analogue to Step 4 Part 1 of the previous algorithm [Praveen2023_EGF].
In order to extend the third part of Step 4 of that same algorithm to the Quasimatrix generalization, we need to establish a valid approximation to the exponential map so that we can map interpolated points in the tangent space back to the Stiefel manifold. The key concept is that of a retraction on a manifold, which may be thought of as a first order approximation to the exponential map. In their previous algorithm Praveen et al.[Praveen2023_EGF] use QR factorization as their approximation to the exponential map, as it is retraction on the finite dimensional Stiefel manifold. Altman et al. [Altmann2022_Stiefel] show that Quasimatrix QR factorization is a retraction on our infinite dimensional Stiefel manifold $S_{K}$, so it follows that we may simply use the Quasimatrix version of QR factorization as our approximation to the exponential map in our updated algorithm.
@@
    $\Gamma^{\mathcal{U}_{0}}_{j}$ $=\mathcal{U}(s;\theta_{j})-\mathcal{U}(s;\theta_{0})\ \text{sym}\left(\mathcal{U}^{*}(s;\theta_{0})\ \mathcal{U}(s;\theta_{j})\right),$
    $\Gamma^{\mathcal{V}_{0}}_{j}$ $=\mathcal{V}(x;\theta_{j})-\mathcal{V}(x;\theta_{0})\ \text{sym}\left(\mathcal{V}^{*}(x;\theta_{0})\ \mathcal{V}(x;\theta_{j})\right),\quad\text{sym}\left(Y\right)\coloneqq(Y+Y^{*})/2.$
@@
2.7.1 Correcting sign and order for singular functions
@@
We observe that the first two singular functions swap at some critical value of $\theta$. The $k^{\text{th}}$ singular value for the Green’s functions of the operator $\mathcal{L}_{\theta}$ are given by $\sigma_{k}=1/(\theta^{2}-(k\pi)^{2})$. If we compare the first two modes, they satisfy $|\sigma_{1}|>|\sigma_{2}|$, when $\theta<\sqrt{5/2}\pi$, and $|\sigma_{1}|\leq|\sigma_{2}|$ otherwise. This implies that the first two left singular functions are swapped when $\theta>\sqrt{5/2}\pi$, as illustrated by Fig. 7, and that the manifold interpolation technique will perform poorly in such cases. The same phenomena occurs for the higher singular functions at larger values of $\theta$. We propose to reorder the singular functions and the associated singular values of the discovered basis at the given interpolation parameter to match the ones at the origin point $\theta_{0}$, where we lift to the tangent space of the infinite-dimensional analogue for Stiefel manifold. For a given parameter, $\theta_{j}$, and mode number, $1\leq k\leq K$, we select the $k^{\text{th}}$ left singular function at $\theta_{j}$ to be the one with minimal angle with the $k^{\text{th}}$ left singular function at parameter $\theta_{0}$, i.e.,
$\mathcal{U}_{k}(s;\theta_{j})=\mathcal{U}_{\ell^{\prime}}(s;\theta_{j}),\quad\text{where}\quad\ell^{\prime}=\text{match}(k),$
$\text{match}(k)=$ $\operatorname*{arg\,max}_{1\leq\ell\leq K,\ell\neq\text{match}(k^{\prime})\forall 1\leq k^{\prime}<k}|\langle\mathcal{U}_{\ell}(s;\theta_{j}),\mathcal{U}_{k}(s;\theta_{0})\rangle|,\quad 1\leq k\leq K.$
where $\langle\cdot,\cdot\rangle$ denotes the inner product in the continuous $L^{2}$ sense. Note that the reordering is done by matching the left singular functions, starting from the left singular function with the highest singular value, without replacement. When the order of the left singular functions has been changed by this procedure, we also re-order the corresponding singular values and right singular functions to preserve the value of the Green’s function.
@@
Figure 7: (A) Magnitude of the first two singular values of the Green’s function associated with the one-dimensional Helmholtz equation. The first two corresponding left singular functions are swapped when the frequency, $\theta$, increases beyond the critical value of $\theta_{\text{crit}}=\sqrt{5/2}\pi\approx 5$, as displayed in (B) and (C).
@@
2.8 A note on computational complexity
The following subsections provide a note on the computational complexity of all the key operations used in the chebgreen package:
[…]
2.8.1 chebfun2—
Constructing a chebfun2 for a bivariate function $g(x,y)$ (Section 2.4) requires sampling the function on an $N\times N$ Chebyshev tensor grid. $N$ depends on the accuracy prescribed in $x$ and $y$ directions. This results in a matrix of sampled values. The process then involves computing an approximate singular value decomposition (SVD) of this matrix. To obtain a near-optimal rank $K$ approximation, the computational complexity is $\mathcal{O}(K^{2}N+K^{3})$ [Townsend_Chebfun2].
2.8.2 QR decomposition
The implementation for a QR decomposition of a column Quasimatrix, $\mathcal{A}$, of size $(\infty\times K)$ (Section 2.5), where $K$ is the rank of the column Quasimatrix, is analogous to the Householder QR decomposition for a conventional matrix when working in a space with an $L^{2}$ inner product. Since the columns of a column Quasimatrix are represented by a chebfun, the additional cost is the conversion between the Chebyshev coefficients, which represent the chebfun and the corresponding values of the function at Chebyshev points. This conversion utilizes the Fast Fourier Transform (FFT) for coefficients to values, and the Inverse Fast Fourier Transform (IFFT) for values to coefficients. Assuming each chebfun in $\mathcal{A}$ is represented by $M$ coefficients, the cost for these transforms is $\mathcal{O}(KMlog(M))$. This cost is dominated by the cost of computing the matrix Householder QR decomposition, which has a complexity $\mathcal{O}(d^{2}D)$, where $d=\min(M,K)$ and $D=\max(M,K)$. Given that generally $K<M$, the total computational complexity for the QR decomposition is $\mathcal{O}(K^{2}M)$.
2.9 Singular Value Expansion
A similar argument is valid for computing the Singular Value Expansion (SVE) of a chebfun2, which is internally stored as a CDR decomposition. Let $C$ be a $(\infty\times K)$ Quasimatrix, $R$ be a $(L\times\infty)$ Quasimatrix, and $D$ be a $(K\times L)$ matrix. Assume $M$ is the number of coefficients used to represent the chebfun-s in both $C$ and $R$. Without loss of generality, assuming $K>L$, the dominant cost in computing the SVE arises from the QR decompositions performed in Algorithm 1: Step 1. Hence, the computational complexity for computing the SVE is $\mathcal{O}(K^{2}M)$.
2.9.1 Interpolation
For interpolating between models of Green’s functions, the primary computational cost lies in lifting the left and right singular functions for all interpolant models to the tangent space at the reference basis (Section 2: Step 1). Assuming that each interpolant model has a rank $K$ singular value expansion and that the chebfun-s representing the left and right singular functions are represented by $M$ coefficients, the computational complexity for interpolating between $N_{\text{models}}$ interpolant models is $\mathcal{O}\left(N_{\text{models}}K^{2}M\right)$.
3 Theoretical error analysis
In this section we consider each of the main sources of error for our interpolation approach: that from the orthogonal projection onto the tangent space of the Hilbert Stiefel manifold $S_{K}$; that from the Lagrange polynomial interpolation of quasimatrices in the tangent space; and, that from the $QR$-based retraction back to the manifold. Although the error analysis for the Lagrange polynomial interpolation is a classical result, little attention has been given to error analysis for the other two steps in general, but especially in the Hilbert manifold setting. Furthermore, no attention has been given to quantifying the size of the region about a base point by which these methods are valid. The proofs for the results presented in this section are given in Appendix B.
3.1 Error due to the projection step
The projection mapping computed in Theorem 1 is the first instance in Algorithm 1 of introducing approximation error after the Green’s functions are represented as chebfuns. Exploiting properties of the Hilbert Stiefel manifold $S_{K}$, we demonstrate that the error in this step decays at least cubically as the distance between points approaches zero in a suitable region of the domain. Before we do this, however, we establish a concrete region within which the Riemannian exponential map is guaranteed to be a local diffeomorphism: this quantity is known as the radius of injectivity.
Lemma 1.
The radius of injectivity of the Riemannian exponential map on the Hilbert Stiefel manifold $S_{K}$ of $H=(L^{2}(\Omega))^{K}$ is $\pi$.
This result tells us that, if $\Phi$ is our base point and $\Psi$ another point of interest, then the the map $\exp_{\Phi}\colon T_{\Phi}S_{K}\rightarrow S_{K}$ is a local diffeomorphism, say on some neighborhood $U$ of $\Phi$ in $T_{\Phi}S_{K}$ whose image is a geoedsic ball of at most radius $\pi$ in the geodesic distance on $S_{K}$. While the geodesic distance may seem to be an abstract way to quantify this region, given that our method does not use this distance directly, however it does furnish an analogous result for the $H$-norm on a tangent vector: that is if $\Gamma=\exp_{\Phi}(\Psi)$, then the quantity $\|\Gamma\|_{H}$ is equal to the geodesic distance between $\Phi$ and $\Psi$ [stoye2024injectivityradiusstiefelmanifold].
Now that we have quantified a region on which the exponential map is a diffeomorphism, we may begin quantifying error estimates, starting with that which is due to the orthogonal projection on $S_{K}$.
Proposition 1 (Cubic approximation error of the projection mapping onto the tangent space).
Let $S_{K}$ be the Hilbert-Stiefel manifold endowed with the Riemannian metric induced from
$H=\bigl{(}L^{2}(\Omega)\bigr{)}^{K},$
and let $\phi\in S_{K}$. Suppose the neighborhood $U$ of the origin in $T_{\phi}S_{K}$ contains only tangent vectors $v$ such that $\|v\|_{H}\ <\pi$. Let $P_{\phi}:H\to T_{\phi}S_{K}$ be the orthogonal projection onto $T_{\phi}S_{K}$ . Then there exists a constant $C>0$ such that, for every $y\in W$ with $y=\exp_{\phi}(v)$, where $W$ is the image of the exponential map on $U$, we have
$\|v-P_{\phi}(y-\phi)\|_{H}\leq C\,\|y-\phi\|_{H}^{3}.$
This is a somewhat surprising result. One might assume that, doing a Taylor expansion in order to compute the error between a projected point and an actual point that error would be quadratic in the distance between points as this distance approaches zero. But it turns out that, even in the Hilbert manifold case [biliotti2006exponentialmapweakriemannian], in a region where the Riemannain exponential map is a local diffeomorphism, the second order term of the Taylor expansion of the exponential map is identically zero [Boumal_2023]. This is the key to obtaining the above result.
3.2 Error due to the $QR$-based retraction
After interpolating tangent vectors in the tangent space, we map back down to the Hilbert Stiefel manifold using a mapping that sends a quasimatrix to the $Q$ in its $QR$ factorization. This step represents the bulk of the error in the entire interpolation process, dominating the error in both the projection and Lagrange polynomial interpolation steps. The following result is well-established in the finite dimensional case, and is a consequence Taylor’s theorem for Banach spaces.
Proposition 2 (Quadratic error for a retraction).
Let $M\subset H$ be an embedded Hilbert manifold in a Hilbert space $H$, so that each tangent space $T_{\phi}M$ is a linear subspace of $H$. Suppose that
$R:TM\to M,\quad(\phi,\eta)\mapsto R(\phi,\eta)$
is a $C^{2}$ retraction, that is that:
  1. (i)
    $R(\phi,0)=\phi$.
  2. (ii)
    $DR(\phi,0)[\eta]=\eta$ for all $\eta\in T_{\phi}M$.
  3. (iii)
    $R(\phi,\cdot)$ is $C^{2}$. Hence, by continuity there exists a constant $L>0$ (depending on $\epsilon$) such that
    $\|D^{2}R(x,0)[\eta,\eta]\|_{H}\leq L\,\|\eta\|_{H}^{2}\quad\text{for all }\eta\in B_{\epsilon}(0).$
Then for every $x\in M$ and every $\eta\in T_{x}M$ with $\|\eta\|_{H}<\epsilon$, there exists $\delta>0$ such that for all $t$ with $|t|<\delta$ we have
$\|R(x,t\eta)-(x+t\eta)\|_{H}\leq L\,t^{2}\,\|\eta\|_{H}^{2}.$
We now must clarify that the $QR$-based retraction is $C^{2}$. In fact, it is smooth.
Lemma 2 (Smoothness of the Gram–Schmidt map on an open subset of $H$).
Let
$H=\left(L^{2}(\Omega)\right)^{K}$
be a Hilbert space and let $M$ be the Hilbert Stiefel manifold of $H$. Define the map
$qf:U\subset H\to M,$
by applying the Gram–Schmidt process, that is, define
$q_{1}(\phi_{1})=\frac{\phi_{1}}{\|\phi_{1}\|_{H}},$
and for $j=2,\dots,K$, recursively set
$p_{j}(\phi_{1},\dots,\phi_{j})=\phi_{j}-\sum_{i=1}^{j-1}\langle\phi_{j},q_{i}(\phi_{1},\dots,\phi_{i})\rangle_{H}\,q_{i}(\phi_{1},\dots,\phi_{i}),$
$q_{j}(\phi_{1},\dots,\phi_{j})=\frac{p_{j}(\phi_{1},\dots,\phi_{j})}{\|p_{j}(\phi_{1},\dots,\phi_{j})\|_{H}}.$
Then, the mapping
$qf(\phi)=\bigl{(}q_{1}(\phi_{1}),\;q_{2}(\phi_{1},\phi_{2}),\;\dots,\;q_{K}(\phi_{1},\dots,\phi_{K})\bigr{)}$
is smooth on the open set
$U=\{\phi\in H:\|\phi_{i}\|_{H}>0\text{ for }i=1,\dots,K,\text{ and }\{\phi_{1},\dots,\phi_{K}\}\text{ is linearly independent}\}.$
It follows that the error associated with the $QR$-based retraction employed in our algorithm decays quadratically as the norm distance between the quasimatrices approaches zero.
3.3 Combined error
Now we consider the total error associated with the interpolation scheme. In the projection step, we obtain cubic order error. After this step, we use Lagrange polynomial interpolation, which has error of order $\mathcal{O}(h^{n+1})$, where $n$ is the number of interpolants and $h$ the step size. Finally, we get quadratic error in the retraction mapping back to the manifold, which dominates the error of the other two steps in the interpolation scheme.